OpenStudy (anonymous):

log6(x+5)+log6(x+4)=2 Can anyone show me how to do this problem?

OpenStudy (anonymous):

\[\log_{6}(x+5)+\log_{6}(x+4)=2 \]\[\log_{6}(x+5)(x+4)=2 \]\[(x+5)(x+4)=6^{2}\](x+5)(x+4)=36\[x ^{2}+9x-16=0\] Solve using quadratic equation.